OpenStudy (anonymous):

I need help solving this problem and showing my work i dont quite understand the steps on how to do it. y=x-1 x+y=7

OpenStudy (anonymous):

Use substitution: substitute the value of y (y = x-1) into the second equation and solve for x: x + (x - 1) = 7 x + x -1 = 7 2x - 1 = 7 +1 2x = 8 /2 x = 4 Now put in the value of x to solve for y: x + y = 7 4 + y = 7 -4 y = 3

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

x+(x-1)=y the (x-1) is because y=x-1 so it replaced y. 2x-1=7 2x=6 x=3 3+y=7 y=5

OpenStudy (anonymous):

@rickybobby13 you solved the first part incorrectly: In order to get the 2x alone in 2x-1=7 you need to add the 1 to both sides.

OpenStudy (anonymous):

@jazy yes thanks so much i got it now.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

my bad guys